OpenStudy (anonymous):

What is an equation of the line, in point-slope form, that passes through the given point and has the given slope? Point: (6,-7) Slope= 4

OpenStudy (anonymous):

Use the point slope equation and plug in the given information to solve for b. y = mx + b -7 = 4(6) + b After you find b, rewrite the above equation using the slope, m, and the value to got for b.

OpenStudy (anonymous):

does that help